a varible z varies jointly with x and the square of y. if z=5 when x=3 and y=4, find the constant of proportionality, k.
i can not figure this out. please help me. thanks so mcuh :)

Answer by stanbon(75887)

a varible z varies jointly with x and the square of y. if z=5 when x=3 and y=4, find the constant of proportionality, k.
The first sentence means z=kxy^2.
You need to solve for "k" as follows:
5=k*3*5^2
5=k(75)
k=1/15
Cheers,
Stan H.